# 3D Hole Filling using Deep Learning Inpainting

### v1

@@ Abstract
The current work presents a novel methodology for completing 3D surfaces produced from 3D digitization technologies in places where there is a scarcity of meaningful geometric data. Incomplete or missing data in these three-dimensional (3D) models can lead to erroneous or flawed renderings, limiting their usefulness in a variety of applications such as visualization, geometric computation, and 3D printing. Conventional surface estimation approaches often produce implausible results, especially when dealing with complex surfaces. To address this issue, we propose a technique that incorporates neural network-based 2D inpainting to effectively reconstruct 3D surfaces. Our customized neural networks were trained on a dataset containing over 1 million curvature images. These images show the curvature of vertices as planar representations in 2D. Furthermore, we used a coarse-to-fine surface deformation technique to improve the accuracy of the reconstructed pictures and assure surface adaptability. This strategy enables the system to learn and generalize patterns from input data, resulting in the development of precise and comprehensive three-dimensional surfaces. Our methodology excels in the shape completion process, effectively filling complex holes in three-dimensional surfaces with a remarkable level of realism and precision.
@@
Figure 1: Comparison of different shape completion techniques. The ground truth mesh with holes (leftmost) is shown alongside various reconstruction methods. Smooth reconstruction using a popular approach [42], Poisson method and Ramesh method (these methods offer the best mesh distance measures for the hole filling task) and our proposed method (rightmost). The proposed method demonstrates superior fidelity in filling complex holes, particularly in regions with intricate details, as seen in the comparison with other techniques.
@@ 1 Introduction
In recent years, there has been a growing prominence of 3D digitization, which has had a significant impact on various disciplines. This technology enables the creation of highly detailed digital replicas of physical objects. Digital models acquired through the use of scanners or photogrammetry are extensively utilized in various fields such as architectural modeling, cultural heritage preservation, reverse engineering, medical simulations, product design, and even in the facilitation of asset creation for the cinema and video game industries. Nevertheless, the efficacy of these 3D digital models is contingent upon the quality of their mesh. The quality in question is not solely associated with precision and resolution, but also encompasses the property of watertightness. An incomplete surface can result in visual artifacts in the final render, hinder the ability to perform geometric computations on the 3D model, or even cause failures in 3D printing.
In parallel to the acquisition methods, the computer graphics community has developed algorithms for model repair. In the specific context of addressing voids on a surface, scholars have employed two primary methodologies: volumetric and surface-oriented algorithms. Volumetric algorithms execute a comprehensive remeshing process on the surface by converting an intermediate point cloud. They ensure the creation of models that are free from gaps or leaks. However, these algorithms often result in a reduction of the surface resolution and a loss of intricate features that were initially captured during the digitization process [6]. In contrast, surface-oriented algorithms operate at a localized level, making modifications to the model solely in the regions directly associated with the hole. Traditional surface-oriented hole-filling algorithms [42, 74] fail to consider the global characteristics of the model or the resemblance to the adjacent surface of the hole. Therefore, a distinct category of algorithms, known as context-based algorithms, can be found in the existing literature. These algorithms aim to replicate the characteristics of the surrounding surface in the vicinity of the area that is missing [29, 71, 17]. In recent years, the increasing prominence of deep learning methods has integrated generative models to accomplish this objective. These models primarily utilize the point cloud representation, thereby disregarding the surface’s topological information [32, 72, 68, 76].
Undoubtedly, deep learning methods have demonstrated remarkable performance in point cloud completion tasks[69, 47, 64], as well as in 2D image completion tasks [84]. The present work introduces a novel approach for filling holes in 3D surface models by leveraging the concept of 2D inpainting and the adaptability of deep learning.
The utilization of a point cloud representation of the object is not relied upon, as a novel patch of polygons is generated to enclose the surface. The consideration of the topology and original shape of the model is achieved through the utilization of a planar representation of the curvature at vertices, which serves as the 2D image that is processed by the neural network. The image containing the holes serves as the input, while the resulting image after applying the trained algorithm is referred to as the inpainted image. The method employed in our study integrates surface-oriented techniques with context-aware information and leverages the creative potential of deep learning. Recently, there has been a utilization of inpainting techniques to tackle the issue of hole-filling in disk-homeomorphic surfaces [51, 25]. These approaches have demonstrated promising outcomes by leveraging a substantial neighborhood surrounding the hole as contextual information for the reconstruction process. In this study, we present a novel approach that extends the existing methods for mesh completion. Our proposed method aims to address the mesh completion task by incorporating information not only from the damaged mesh itself, but also from other polygonal models that share similarities with the one requiring repair.
Our primary goal has been to improve the precision of the reconstructed surface and to fully automate the restoration process, thereby surpassing the constraints of general-purpose inpainting neural networks, as those used in similar works [33]. In order to accomplish these objectives, we have trained the inpainting networks with over 1 million curvature images from a large 3D model dataset. The use of ad-hoc trained neural networks improves the accuracy of the deep learning inpainting algorithm, as they are specialized in a well-defined type of images with the same color range. Then, a coarse-to-fine surface deformation algorithm is implemented to conform the newly generated surface patch to the anticipated outcome. Additionally, we have enhanced this autonomous virtual sculpting algorithm to generate surfaces that are more detailed and precise. Our method is noteworthy for its ability to optimize the repair of 3D models with a substantial number of polygons and intricate shapes, as exemplified in Fig. 1, transcending the limitations of conventional methods that typically address simpler models.
The remainder of the paper is organized as follows. The purpose of Section 2 is to contextualize our work and provide the most recent and pertinent literature that motivates us in our research. Subsequently, Section 3 illustrates the dataset generation, network training, and surface reconstruction processes. We conduct a quantitative and qualitative comparison of our reconstructed surfaces with some of the most advanced algorithms in Section 4. Finally, in Section 5, we discuss the current limitations and challenges that will inspire our future research, as well as illustrate a few interesting implications of this approach.
@@ 2.3 Shape Completion using Inpainting Techniques
3 Methodology
@@
The challenge of restoring complex objects with multiple holes is tackled in this work through the utilization of a curvature-based approach. This approach entails the creation of a curvature feature image and the subsequent application of inpainting techniques to effectively fill in the missing components. The iterative process of deformations ensures the successful restoration of complex surfaces by inferring missing geometry over the mesh surface until alignment with the inpainted curvature image is achieved. The focus of our research is on automating and improving the methodology used to address the limitations of generic inpainting networks.
This procedure is further clarified upon in this section. At the initial phase, we generate a dataset of 3D models that are all of a specific dimension and manifoldness. The image dataset that will be employed in the deep learning training is built upon this curated collection of meshes. The entire process entails the conversion of the 3D mesh representations into 2D images, thereby guaranteeing that the resultant images accurately depict the geometrical details of the original surface. These 3D meshes will not only be used to produce 2D images but also as a benchmark for assessing the efficacy of our shape completion algorithm. Section 3.1 delineates all aspects of the procedure.
Following that, we proceed to train a range of cutting-edge ad-hoc architectures for image inpainting utilizing our image dataset, as depicted in Section 3.2. The automatic algorithm that conducts virtual sculpting of the mesh surface will be guided by the resulting repaired images. This method, as illustrated in 3.3, offers a comprehensible process that results in an accurate solution to the 3D model shape reconstruction problem.
@@ 3.1 Creation of the training dataset
As outlined in 3.1.3, the complete models undergo a secondary procedure where their 3D surfaces are partitioned into patches and subsequently parametrized to yield multiple 2D surfaces per mesh. Subsequently, artificial holes are introduced into the images, and through the implementation of traditional augmentation techniques, a dataset consisting of over one million images is obtained. This dataset is then utilized for training the inpainting neural network.
@@ 3.1.1 Selection of 3D models
To ensure the success of the parametrization or reconstruction process, it is imperative that all meshes in the dataset possess the quality of being watertight, meaning they should not contain any holes. Additionally, they should be manifold, indicating that they are continuous and have no self-intersections. Furthermore, the meshes should consist of a single component with a substantial number of polygons. The aforementioned characteristics play a pivotal role, as the correct topology enables us to execute geometric operations on the mesh surface while repairing the affected areas, thereby ensuring the efficacy of our reconstruction algorithm.
@@ 3.1.2 Creating holes in 3D models for testing
Figure 3: Systematic approach to perforation simulation on 3D manifold models. Initial 3D mesh model without perforations and selection of starting vertices using a farthest point sampling strategy.
@@ 3.1.3 Curvature images for training
3.2 Neural Network Training
@@
Figure 8: Scheme of our training process. Every approach selected is fine-tuned using our image dataset and then the results are evaluated by comparing with the ground truth images.
@@
3.3 3D Reconstruction Algorithm
@@
Next, the mesh is segmented into local patches around every hole to preserve contextual information. The curvature values are used to color the per-vextex of these patches, which are parameterized to produce planar representations of the surface. The hole mask image is the result of the newly generated vertices in the previous phase.
@@
Figure 10: Deformation process. The curvature color calculation zone outlines the process of identifying color variations between the area where the coarse hole reconstruction takes place and the corresponding curvature inpainted image. The deformation process zone encompasses the two main aspects of deformation: the ROI, which is the coarse reconstructed patch, and a set of control points that exhibit noticeable color variations. The vertices are shifted based on the vertex displacement scheme, as the deformation requires moving them in different directions relative to the normal. Finally, we present the reconstruction following the deformation process, showcasing the updated curvature colors and the final reconstructed surface.
@@
4 Results
Our approach has been assessed quantitatively and qualitatively through a series of experiments. In order to guarantee impartial and equitable comparisons and measurements, we offer a concise summary of the experimental conditions in Section 4.1. Consequently, in Section 4.2, we examine the quantitative experiments and present the corresponding results. Lastly, we investigate the qualitative experiments in Section 4.3 to illustrate the quality of the reconstructed meshes and the performance of our approach.
@@ 4.3 Quantitative Results
In our comprehensive assessment of the methodology, we concentrated on the comparison of the Hausdorff distance error between the original mesh and the repaired mesh. Table 4 contains the summarized results of this comparison of the four models that were previously employed. These metrics will further facilitate the understanding of the visual reparation results that were previously provided.
Our method consistently demonstrates a significantly reduced maximum distance in comparison to alternative approaches when the Hausdorff distance is taken into account. This finding implies that our methodology is proficient in mitigating substantial errors within the reconstructed meshes. In addition, the mean distance metric consistently outperforms other methods, suggesting that our reconstructions produce more consistent and stable results across the test models.
@@
We conducted a comprehensive evaluation of the repairing algorithms on the entire mesh dataset to verify that our approach substantially outperforms others in terms of Hausdorff distance. The experiment conditions outlined in Section 4.1 were taken into account during the execution of these operations.
@@
5 Conclusions and future works
Based on the results presented in the previous section, our approach excels in mesh repair tasks for a variety of reasons. Firstly, by utilizing deep learning techniques, our method is able to acquire intricate surface curvature color patterns, leading to repairs that are highly accurate. Additionally, our thorough repair process focuses on preserving the original mesh shape and details. This results in repaired surfaces that closely resemble the ground truth meshes, without any over-smoothing or unnatural deformations that are often found in other methods. Our quantitative evaluations provide evidence of this. Our approach brings unique strengths that greatly contribute to the field of shape completion tasks.
Although our approach has shown significant strengths in mesh repair tasks, it is crucial to recognize certain limitations. The effectiveness of our method can be affected by the quality and diversity of the training dataset due to the use of deep learning techniques and extensive training data. Inadequate or skewed training data may lead to less than optimal repairs. In addition, handling severely damaged or incomplete meshes can pose challenges when trying to preserve mesh details during the repair process. A noteworthy example is when meshes show the existence of separate elements or faces in the geometry. Our method might encounter difficulties in effectively repairing such meshes, as the segmentation and curvature calculations involved necessitate watertight geometries.

### v2

# SR-CurvANN: Advancing 3D Surface Reconstruction through Curvature-Aware Neural Networks

@@ Abstract
Incomplete or missing data in three-dimensional (3D) models can lead to erroneous or flawed renderings, limiting their usefulness in applications such as visualization, geometric computation, and 3D printing. Conventional surface-repair techniques often fail to infer complex geometric details in missing areas. Neural networks successfully address hole-filling tasks in 2D images using inpainting techniques. The combination of surface reconstruction algorithms, guided by the model’s curvature properties and the creativity of neural networks in the inpainting processes should provide realistic results in the hole completion task. In this paper, we propose a novel method entitled SR-CurvANN (Surface Reconstruction Based on Curvature-Aware Neural Networks) that incorporates neural network-based 2D inpainting to effectively reconstruct 3D surfaces. We train the neural networks with images that represent planar representations of the curvature at vertices of hundreds of 3D models. Once the missing areas have been inferred, a coarse-to-fine surface deformation process ensures that the surface fits the reconstructed curvature image. Our proposal makes it possible to learn and generalize patterns from a wide variety of training 3D models, generating comprehensive inpainted curvature images and surfaces. Experiments conducted on 959 models with several holes have demonstrated that SR-CurvANN excels in the shape completion process, filling holes with a remarkable level of realism and precision.
@@
Figure 1: Comparison of different shape completion techniques. The ground truth mesh with holes (leftmost) is shown alongside various reconstruction methods. Smooth reconstruction using a popular approach [41], Poisson method [37] and Ramesh method [9] (these methods offer the best mesh distance measures for the hole filling task) and SR-CurvANN (rightmost). The proposed method demonstrates superior fidelity in filling complex holes, particularly in regions with intricate details, as seen in the comparison with other techniques.
@@ 1 Introduction
In recent years, 3D digitization has grown in popularity, affecting a wide range of fields. This technology enables detailed digital replicas of actual objects. Architectural modeling, cultural heritage preservation, medical simulations, and the development of film and video game assets all make use of scanned or photogrammetric digital models. However, mesh quality is what determines the efficacy of a 3D digital model. Uncomplete surfaces can cause render artifacts, geometric computation mistakes, and 3D printing problems.
To deal with surface imperfections, researchers use both volumetric and surface-oriented approaches. To assure hole-free models, volumetric approaches convert the mesh to an intermediate point cloud, but may degrade surface resolution and lose sensitive details [6]. Surface-oriented algorithms [41, 73] prioritize localized modifications near the hole above global model attributes. Context-based approaches simulate the surrounding surface of the missing area [28, 70, 17].
More recently, deep learning methods have facilitated the use of generative models to tackle surface completeness. These models mostly employ point cloud representations and ignore topological information [31, 71, 67, 75]. They demonstrated outstanding performance in point cloud [68, 46, 63] and 2D image [83] completion tasks. Convolutional neural networks (CNNs) have also demonstrated to effectively help in repairing holes by leveraging contextual information from the surrounding neighborhood in a 2D domain [25, 44].
Inspired by these recent advances, we propose a novel method called SR-CurvANN (Surface Reconstruction through Curvature-Aware Neural Networks). This method integrates surface-oriented techniques with curvature-aware information and leverages the creative potential of deep learning processes to address the mesh completion task. Firstly, we generate a dataset of more than one million curvature images from an extensive collection of 3D models, and then apply random hole masks to these images. Using this image database, we train inpainting neural networks to repair holes in a concrete color palette spectrum. Once trained, these ad-hoc neural networks can infer missing parts in curvature images with real holes, thereby parametrizing and repairing the patches of the incomplete mesh. Finally, we apply an automatic curvature-guided surface reconstruction algorithm to the freshly created surface patch until it aligns accurately with the expected result. By following these steps, we recreate 3D surfaces with a remarkable level of realism and accuracy, especially in areas with intricate geometry. Therefore, the primary contributions of this work might be summarized as follows:
  • We propose a fully automated method, named SR-CurvANN, which integrates image inpainting and virtual sculpting to reconstruct surfaces with a substantial number of polygons and intricate shapes, surpassing the limitations of existing methods that typically address simpler models.
  • We show that ad-hoc trained neural networks outperform general-purpose inpainting networks on the dataset of curvature images, leading to a more accurate and precise surface completion.
  • We have defined a process to create a vast collection of training images from authentic 3D models featuring holes.
@@ 2.3 Shape Completion using Inpainting Techniques
3 Overview
Figure 2: Overview of SR-CurvANN.1) Creation of curvatures images and holes dataset. We use a huge collection of 3D models to make a dataset of more than a million images of curvature, adding random hole masks to these images. 2) Training curvature-aware inpainting NNs. We use this database of images to teach inpainting artificial neural networks how to fix holes in a specific color palette spectrum. 3) Inpainting over masked curvature image. Once they are taught, these ad-hoc neural networks can infer missing parts of curvature images, prompting requests for them to inpaint images extracted from 3D meshes previously unused. 4) Deformation Algorithm. An automatic curvature-guided surface reconstruction algoritm is applied on the newly made surface patch until it matches the desired outcome perfectly.
This article introduces a method dubbed SR-CurvANN (Surface Reconstruction using Curvature-Aware Neural Networks) that integrates surface-based approaches and curvature-sensitive data to train deep learning algorithms for predicting the shape of the missing portion of the mesh.
As illustrated in the visual diagram of SR-CurvANN shown in Figure 2, our method may be described into four main stages:
  1. 1.
    Data preparation: We extract over 1 million curvature images from a large collection of topologically accurate 3D models. The images depict the three-dimensional curvature at the vertices, shown in a two-dimensional format using a color-coded scale. In addition, each curvature image is associated with a collection of randomly generated hole masks, which enhances the diversity of the training dataset.
  2. 2.
    Training of neural networks: To train the inpainting neural networks, we use the database of curvature images and their corresponding hole masks as input. Such ad-hoc trained neural networks improve the precision of the deep learning inpainting process because they are specifically built to repair images with a well-defined color range.
  3. 3.
    Inference over real damaged models: The curvature-aware neural network receives a 2D curvature-based image from a 3D surface patch that contains a hole and is able to effectively reconstruct the missing area. The neural network infers the curvature of the missing part from the patch itself, as well as from features from other similar polygonal models, thanks to prior training.
  4. 4.
    Coarse-to-Fine surface deformation algorithm: Upon completion of the inference phase, we apply a coarse-to-fine surface deformation algorithm on the 3D mesh. This algorithm aligns an initially simple surface patch with the expected result, guaranteeing that the reconstructed surface is comparable to the real geometry and topology of the model. As Figure 1 shows, this stage enables the final patch to achieve an extraordinary degree of accuracy and detail.
4 Method
@@
Our SR-CurvANN method entails the creation of a curvature feature image and the subsequent application of inpainting techniques to effectively fill in the missing components. The focus of our research is on automating and improving the methodology used to address the limitations of generic inpainting networks.
In the following, we describe in Section 4.1 the selection of the 3D models that are used for the training images dataset. The entire process entails the conversion of the 3D mesh representations into 2D images, thereby guaranteeing that the resultant images accurately depict the geometrical details of the original surfaces. We will not only use these 3D meshes to produce 2D training images, but also use them as a benchmark to evaluate the effectiveness of our shape completion algorithm. Following that, in Section 4.2 we explain the second stage of our method, which is implemented using different cutting-edge ad-hoc architectures for image inpainting to benchmark their accuracy. Once trained, the different inpainting approaches are subjected to evaluation, requesting them to complete 3D meshes with missing parts and comparing their performance, as depicted in Section 4.3. The automatic algorithm that conducts virtual sculpting of the mesh surface is explained in Section 4.4. This method, as will be explained in Section 5, results in an accurate solution to the 3D model shape reconstruction problem.
@@ 4.1 Creation of the training dataset
As outlined in 4.1.3, the complete models undergo a secondary procedure where their 3D surfaces are partitioned into patches and subsequently parametrized to yield multiple 2D surfaces per mesh. Subsequently, artificial holes are introduced into the images, and through the implementation of traditional augmentation techniques, a dataset consisting of over one million images is obtained. This dataset is then utilized for training the inpainting neural network.
@@ 4.1.1 Selection of 3D models
To ensure the success of the parametrization or reconstruction process, it is imperative that all meshes in the dataset possess the quality of being watertight, meaning they should not contain any holes. Additionally, they should be manifold, indicating that they are continuous and have no self-intersections. Furthermore, the meshes should consist of a single component with a substantial number of polygons. The aforementioned characteristics play a pivotal role, as the correct topology enables us to execute geometric operations on the mesh surface while repairing the affected areas, thereby ensuring the efficacy of our reconstruction method.
@@ 4.1.2 Creating holes in 3D models for testing
Figure 4: The initial 3D mesh model follows a farthest point sampling strategy, selecting the starting vertices for the perforations to ensure that each vertex is maximally distant from other vertices in the surface.
@@ 4.1.3 Curvature images for training
4.2 Tranining of Neural Networks
@@
Figure 9: Scheme of SR-CurvANN training process. Every approach selected is fine-tuned using our image dataset and then the results are evaluated by comparing with the ground truth images.
4.3 Inference over curvature images
@@
4.4 Coarse-to-Fine Surface Deformation Algorithm
@@
Next, the mesh is segmented into local patches around every hole to preserve contextual information. The curvature values are used to color the per-vertex of these patches, which are parameterized to produce planar representations of the surface. The hole mask image is the result of the newly generated vertices in the previous phase.
@@
Figure 11: SR-CurvANN Deformation process. The curvature color calculation zone outlines the process of identifying color variations between the area where the coarse hole reconstruction takes place and the corresponding curvature inpainted image. The deformation process zone encompasses the two main aspects of deformation: the ROI, which is the coarse reconstructed patch, and a set of control points that exhibit noticeable color variations. The vertices are shifted based on the vertex displacement scheme, as the deformation requires moving them in different directions relative to the normal. Finally, we present the reconstruction following the deformation process, showcasing the updated curvature colors and the final reconstructed surface.
@@
5 Experiments and evaluation
SR-CurvANN has been assessed quantitatively and qualitatively through a series of experiments. In order to guarantee impartial and equitable comparisons and measurements, we offer a concise summary of the experimental conditions in Section 5.1. Consequently, in Section 5.2, we examine the qualitative experiments and present the corresponding results. Lastly, we investigate the quantitative experiments in Section 5.3 to illustrate the quality of the reconstructed meshes and the performance of our approach.
@@ 5.3 Quantitative Results
In our comprehensive assessment of the method, we concentrated on the comparison of the Hausdorff distance error between the original mesh and the repaired mesh. Table 4 contains the summarized results of this comparison of the four models that were previously employed. These metrics will further facilitate the understanding of the visual reparation results that were previously provided.
Our method consistently demonstrates a significantly reduced maximum distance in comparison to alternative approaches when the Hausdorff distance is taken into account. This finding implies that our method is proficient in mitigating substantial errors within the reconstructed meshes. In addition, the mean distance metric consistently outperforms other methods, suggesting that our reconstructions produce more consistent and stable results across the test models.
@@
We conducted a comprehensive evaluation of the repairing algorithms on the entire mesh dataset to verify that our method substantially outperforms others in terms of Hausdorff distance. The experiment conditions outlined in Section 5.1 were taken into account during the execution of these operations.
@@
6 Conclusions and future works
Based on the results presented in the previous section, SR-CurvANN excels in mesh repair tasks for a variety of reasons. Firstly, by utilizing deep learning techniques, our method is able to acquire intricate surface curvature color patterns, leading to repairs that are highly accurate. Additionally, our thorough repair process focuses on preserving the original mesh shape and details. This results in repaired surfaces that closely resemble the ground truth meshes, without any over-smoothing or unnatural deformations that are often found in other methods. Our quantitative evaluations provide evidence of this. Our method brings unique strengths that greatly contribute to the field of shape completion tasks.
Although SR-CurvANN has shown significant strengths in mesh repair tasks, it is crucial to recognize certain limitations. The effectiveness can be affected by the quality and diversity of the training dataset due to the use of deep learning techniques and extensive training data. Inadequate or skewed training data may lead to less than optimal repairs. In addition, handling severely damaged or incomplete meshes can pose challenges when trying to preserve mesh details during the repair process. A noteworthy example is when meshes show the existence of separate elements or faces in the geometry. We might encounter difficulties in effectively repairing such meshes, as the segmentation and curvature calculations involved necessitate watertight geometries.